OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 5) - 14 = 2(x - 7) + 3 18 25 13 11

OpenStudy (anonymous):

Distribute 3x-15 - 14 = 2x-14+3 Combine like terms 3x-29 = 2x-11 Add 29 to both sides 3x= 2x+18 Subtract 2x from both sides x=18